# Advancing Graph Generation through Beta Diffusion

### v1

@@ 2.2 Forward and reverse beta diffusion processes
2.3 Training GBD
The overall training procedure of GBD is described in Algorithms 3 and 4. We employ the objective function proposed by beta diffusion (Zhou et al., 2023), specifically,
In Equation 10, the KL divergence is evaluated between the following distributions: $q({\mathbf{G}}_{\tau}\,\mid\,\hat{G}_{\theta}({\mathbf{G}}_{t},t))$ is $\mathrm{Beta}(\eta\alpha_{t}\hat{G}_{\theta}({\mathbf{G}}_{t},t),\eta(1-\alpha_{t}\hat{G}_{\theta}({\mathbf{G}}_{t},t)))$, and $q({\mathbf{G}}_{\tau}\,\mid\,{\mathbf{G}}_{0})$ is the same as $q({\mathbf{G}}_{t}\,\mid\,{\mathbf{G}}_{0})$ in distribution. The subscript $\tau$ is introduced to represent a generic graph sample other than ${\mathbf{G}}_{t}$ that is also obtained at time $t$ from the forward diffusion process. The core principle behind the loss function terms can be described as follows: $\mathcal{L}_{\mathrm{sampling}}$ drives the empirical ancestral sampling distribution towards the destination-conditional posterior distribution, while $\mathcal{L}_{\mathrm{correction}}$ corrects the bias on marginal distribution at each time stamp accumulated through the ancestral sampling. These two types of loss terms collectively reduce the divergence between the empirical joint distribution on two graphs sampled from adjacent time stamps in the reverse process, and their joint distribution derived from the forward diffusion process. A positive weight $\omega$ is introduced to balance the effects of these two types of loss terms. We set it to 0.01, following Zhou et al. (2023), and found that this configuration is sufficient to produce graphs that closely resemble the reference graphs without further tuning. To better elucidate the optimization objective, we list out the following analytical expressions:
$=\ln\Gamma(\eta(\alpha_{t-1}-\alpha_{t}){\mathbf{G}}_{0})+\ln\Gamma(\eta-\eta\alpha_{t-1}{\mathbf{G}}_{0})+\ln\Gamma(\eta-\eta\alpha_{t}{\mathbf{G}}_{0})$
$-\ln\Gamma(\eta(\alpha_{t-1}-\alpha_{t})\hat{{\mathbf{G}}}_{0})-\ln\Gamma(\eta-\eta\alpha_{t-1}\hat{{\mathbf{G}}}_{0})-\ln\Gamma(\eta-\eta\alpha_{t}\hat{{\mathbf{G}}}_{0})$
$+\eta\alpha_{t}({\mathbf{G}}_{0}-\hat{{\mathbf{G}}}_{0})\cdot\psi(\eta-\eta\alpha_{t}\hat{{\mathbf{G}}}_{0}),$
$+\eta\alpha_{t}(\hat{{\mathbf{G}}}_{0}-{\mathbf{G}}_{0})\cdot\left(\psi(\eta\alpha_{t}\hat{{\mathbf{G}}}_{0})-\psi(\eta-\eta\alpha_{t}\hat{{\mathbf{G}}}_{0})\right),$
where $\hat{{\mathbf{G}}}_{0}:=\hat{G}_{\theta}({\mathbf{G}}_{t},t)$.
Here we reference the $d_{\phi}$ notation of Banerjee et al. (2005) to represent the Bregman divergence. As stated in Lemmas 3-5 of Zhou et al. (2023), one can apply Proposition 1 of Banerjee et al. (2005) to show that both $\mathcal{L}_{\mathrm{sampling}}$ and $\mathcal{L}_{\mathrm{correction}}$ yield the same optimal solution that legitimates the usage of $\hat{{\mathbf{G}}}_{0}$ in the reverse diffusion process.
Property 1.
@@ 2.3 Training GBD
2.4 Exploring the design space of GBD
Data transformation.
Concentration modulation.
Figure 2: Reverse Process on Graph Edges: Each edge $(u,v)$ undergoes a reverse process with specific $\eta_{u,v}$ parameters, influenced by $\text{deg}(u)$ and $\text{deg}(v)$. The colorbar indicates the normalized degree of nodes in the final graph, with brighter colors (shifting towards yellow) denoting higher degrees. Nodes in the adjacency matrices shown in the bottom row are reordered based on decreasing node degree.
Another hyperparameter that offers a more refined adjustment to the mixing rate is the concentration parameter $\eta$. Higher values of $\eta$ reduce the variance of the fractional multipliers ${\mathbf{P}}_{t}$ sampled from their corresponding beta distributions, thus delaying the arrival of the mixing state. Leveraging this property, we have devised a simple yet effective modulation strategy to differentiate the mixing time for various graph substructures.
Specifically, we assign higher $\eta$ values to “important positions” within a graph, such as edges connecting high-degree nodes or edges redeemed as significant based on domain knowledge, such as the carbon-carbon bond in chemical molecules. For instance, when modulating $\eta$ from node degrees, the exact operation executed upon the $\eta$ values for edge $(u,v)$ and for the features of node $u$ can be mathematically expressed as
Here we first prepare several levels of $\eta$ values, then utilize two assignment functions, namely $g_{A}(\cdot)$ and $g_{X}(\cdot)$, to map the node degrees (or their percentile in the degree population) to one of the choices of the $\eta$ values. We have observed that this operation indeed prolongs the presence of these substructures during the forward diffusion process, which in turn leads to their earlier emergence compared to the rest of the graph during the reverse process.
We visualize the reverse process from two perspectives in Figure 2. We first obtain the $\eta_{u,v}$ by degrees retrieved from the training set before sampling and then generate graph through reverse beta diffusion. From the top row, we observe that edges linked to nodes with higher degrees (indicated by brighter colors) appear first, followed by other edges. From the bottom row, it is evident that edges connected to the first five nodes, which have higher degrees, are identified first and then progressively in descending order of degree. Notably, the nodes of the adjacency matrices in the bottom row are reordered by decreasing node degree of the final graph. Additionally, we can also find the predicted graph of GBD converges in an early stage to the correct topology. We attribute the enhanced quality of generated graphs to the early emergence of these “important substructures,” which potentially enhances the reliability of generating realistic graph structures. Furthermore, we find this approach particularly appealing because it allows for the flexible integration of graph inductive biases within the diffusion model framework.
Logit domain computation.
Neural-network precondition.
A drawback of the logit-domain computation is its introduction of variability to model training, we address this issue by employing neural-network precondition techniques (Karras et al., 2022). Specifically, precondition involves standardizing ${\mathbf{G}}_{t}$ before passing them to the prediction network $\hat{G}_{\theta}(\cdot)$. In other words, we modify Equation 7 as
$p_{\theta}({\mathbf{G}}_{t-1}\,\mid\,{\mathbf{G}}_{t}):=q({\mathbf{G}}_{t-1}\,\mid\,{\mathbf{G}}_{t},\hat{G}_{\theta}(\tilde{{\mathbf{G}}}_{t},t)),~{}\tilde{{\mathbf{G}}}_{t}=\frac{{\mathbf{G}}_{t}-\mathbb{E}[g_{t}]}{\sqrt{\mathrm{Var}[g_{t}]}}~{}\text{or}~{}\frac{\mathrm{logit}({\mathbf{G}}_{t})-\mathbb{E}[\mathrm{logit}(g_{t})]}{\sqrt{\mathrm{Var}[\mathrm{logit}(g_{t})]}},$ (18)
where $g_{t}$ denotes the elements within ${\mathbf{G}}_{t}$. Based upon the law of total expectation and the law of total variance, Zhou et al. (2023) have derived the following attribute of beta diffusion
Attribute 1.
Given that $g_{t}|g_{0}\sim\mathrm{Beta}(\eta\alpha_{t}g_{0},\eta(1-\alpha_{t}g_{0}))$, one can derive that $\mathbb{E}[g_{t}|g_{0}]=\alpha_{t}g_{0}$ and $\mathrm{Var}[g_{t}|g_{0}]=\frac{\alpha_{t}g_{0}(1-\alpha_{t}g_{0})}{\eta+1}$. Representing $\mathbb{E}[g_{0}]$ and $\mathrm{Var}[g_{0}]$ with $\mu$ and $\sigma^{2}$, we have
$\mathbb{E}[g_{t}]=\alpha_{t}\mu,~{}\mathrm{Var}[g_{t}]=\frac{\alpha_{t}\mu-\alpha_{t}^{2}(\mu^{2}+\sigma^{2})}{\eta+1}+\alpha^{2}_{t}\sigma^{2}.$
Their counterparts in the logit domain are expressed as
$\mathbb{E}[\mathrm{logit}(g_{t})]$ $=\mathbb{E}[\psi(\eta\alpha_{t}g_{0})]-\mathbb{E}[\psi(\eta-\eta\alpha_{t}g_{0})],$
$\mathrm{Var}[\mathrm{logit}(g_{t})]$ $=\mathbb{E}[\psi^{(1)}(\eta\alpha_{t}g_{0})]+\mathbb{E}[\psi^{(1)}(\eta(1-\alpha_{t}g_{0}))]+\mathrm{Var}[\psi(\eta\alpha_{t}g_{0})]+\mathrm{Var}[\psi(\eta(1-\alpha_{t}g_{0}))],$
In addition to Attribute 1, the specific values for $\mathbb{E}[g_{t}]$, $\mathrm{Var}[g_{t}]$, $\mathbb{E}[\mathrm{logit}(g_{t})]$ and $\mathrm{Var}[\mathrm{logit}(g_{t})]$ are also reliant on the distribution of the datum in its initial state. Recall that all variables within ${\mathbf{G}}_{0}$ have been preprocessed using either dummy encoding or CDF transformation. Consequently, the processed variables follow either a discrete distribution or a uniform distribution. With this context, if the original datum follows a discrete distribution, we present the following conclusion.
Remark 1.
If $g_{0}$ has two potential outcomes $\{g_{\min},g_{\max}\}$ with $P(g_{0}=g_{\max})=p$, then
$\mathbb{E}[g_{t}]$ $=\alpha_{t}\left(p\cdot g_{\max}+(1-p)\cdot g_{\min}\right),$
$\mathrm{Var}[g_{t}]$ $=\frac{1}{\eta+1}\left(\mathbb{E}[g_{t}]-\mathbb{E}[g_{t}]^{2}\right)+\frac{\eta}{\eta+1}\left(\alpha_{t}^{2}(p(1-p))(g_{\max}-g_{\min})^{2}\right).$ (19)
The calculation of $\mathbb{E}[\mathrm{logit}(g_{t})]$ and $\mathrm{Var}[\mathrm{logit}(g_{t})]$ relies on a series of components, namely
$\mathbb{E}[\psi(\eta\alpha_{t}g_{0})]$ $=p\cdot\psi(\eta\alpha_{t}g_{\max})+(1-p)\cdot\psi(\eta\alpha_{t}g_{\min}),$
$\mathbb{E}[\psi(\eta-\eta\alpha_{t}g_{0})]$ $=p\cdot\psi(\eta-\eta\alpha_{t}g_{\max})+(1-p)\cdot\psi(\eta-\eta\alpha_{t}g_{\min}),$
$\mathrm{Var}[\psi(\eta\alpha_{t}g_{0})]$ $=p(1-p)\left(\psi(\eta\alpha_{t}g_{\max})-\psi(\eta\alpha_{t}g_{\min})\right)^{2},$
$\mathrm{Var}[\psi(\eta-\eta\alpha_{t}g_{0})]$ $=p(1-p)\left(\psi(\eta-\eta\alpha_{t}g_{\max})-\psi(\eta-\eta\alpha_{t}g_{\min})\right)^{2},$
$\mathbb{E}[\psi^{(1)}(\eta\alpha_{t}g_{0})]$ $=p\cdot\psi^{(1)}(\eta\alpha_{t}g_{\max})+(1-p)\cdot\psi^{(1)}(\eta\alpha_{t}g_{\min}),$
$\mathbb{E}[\psi^{(1)}(\eta-\eta\alpha_{t}g_{0})]$ $=p\cdot\psi^{(1)}(\eta-\eta\alpha_{t}g_{\max})+(1-p)\cdot\psi^{(1)}(\eta-\eta\alpha_{t}g_{\min}),$ (20)
with which one can easily deduce the expressions of $\mathbb{E}[\mathrm{logit}(g_{t})]$ and $\mathrm{Var}[\mathrm{logit}(g_{t})]$ following Attribute 1.
Alternatively, if the original datum follows a uniform distribution upon the support $[g_{\min},g_{\max}]$, then as demonstrated in Zhou et al. (2023), we have
Remark 2.
If $g_{0}$ follows the uniform distribution $\mathrm{Unif}[g_{\min},g_{\max}]$, then
$\mathbb{E}[g_{t}]$ $=\frac{1}{2}\alpha_{t}(g_{\min}+g_{\max}),$
$\mathrm{Var}[g_{t}]$ $=\frac{1}{\eta+1}\left(\mathbb{E}[g_{t}]-\mathbb{E}[g_{t}]^{2}\right)+\frac{\eta}{12(\eta+1)}\left(\alpha_{t}^{2}(g_{\max}-g_{\min})^{2}\right).$ (21)
We denote by $K$ the number of sub-intervals used in the numerical integration based on the Trapezoidal rule. Similar to Remark 1, on the logit domain, we list out the expressions for the components as
$\mathbb{E}[\psi(\eta\alpha_{t}g_{0})]$ $=\frac{1}{\eta\alpha_{t}(g_{\max}-g_{\min})}\left(\ln\Gamma(\eta\alpha_{t}g_{\max})-\ln\Gamma(\eta\alpha_{t}g_{\min})\right),$
$\mathbb{E}[\psi(\eta-\eta\alpha_{t}g_{0})]$ $=\frac{1}{\eta\alpha_{t}(g_{\max}-g_{\min})}\left(\ln\Gamma(\eta-\eta\alpha_{t}g_{\min})-\ln\Gamma(\eta-\eta\alpha_{t}g_{\max})\right),$
$\mathrm{Var}[\psi(\eta\alpha_{t}g_{0})]$ $\approx\max\left(\frac{1}{K}\sum_{i=0}^{K}\frac{\psi^{2}\left(\eta\alpha_{t}\left(g_{\min}+\frac{i}{K}(g_{\max}-g_{\min})\right)\right)}{2^{\delta(i=0)+\delta(i=K)}}-\mathbb{E}[\psi(\eta\alpha_{t}g_{0})]^{2},0\right),$
$\mathrm{Var}[\psi(\eta-\eta\alpha_{t}g_{0})]$ $\approx\max\left(\frac{1}{K}\sum_{i=0}^{K}\frac{\psi^{2}\left(\eta-\eta\alpha_{t}\left(g_{\min}+\frac{i}{K}(g_{\max}-g_{\min})\right)\right)}{2^{\delta(i=0)+\delta(i=K)}}-\mathbb{E}[\psi(\eta-\eta\alpha_{t}g_{0})]^{2},0\right),$
$\mathbb{E}[\psi^{(1)}(\eta\alpha_{t}g_{0})]$ $=\frac{1}{\eta\alpha_{t}(g_{\max}-g_{\min})}\left(\psi(\eta\alpha_{t}g_{\max})-\psi(\eta\alpha_{t}g_{\min})\right),$
$\mathbb{E}[\psi^{(1)}(\eta-\eta\alpha_{t}g_{0})]$ $=\frac{1}{\eta\alpha_{t}(g_{\max}-g_{\min})}\left(\psi(\eta-\eta\alpha_{t}g_{\min})-\psi(\eta-\eta\alpha_{t}g_{\max})\right).$ (22)
Utilizing Attribute 1, one can derive the mathematical expressions that we use to calculate $\mathbb{E}[\mathrm{logit}(g_{t})]$ and $\mathrm{Var}[\mathrm{logit}(g_{t})]$.
3 Related Work
Graph generative models.
Early attempts at modeling graph distributions trace back to the Erdős–Rényi random graph model (ERDdS and R&wi, 1959; Erdős et al., 1960), from which a plethora of graph generative models have emerged. These models employ diverse approaches to establish the data generative process and devise optimization objectives, which in turn have significantly expanded the flexibility in modeling the distribution of graph data. Stochastic blockmodels (Holland et al., 1983; Lee and Wilkinson, 2019), more advanced latent variable models (Airoldi et al., 2009; Zhou, 2015; Caron and Fox, 2017), and their variational-autoencoder-based successors (Kipf and Welling, 2016; Hasanzadeh et al., 2019; Mehta et al., 2019; He et al., 2022) assume that edges are formed through independent pairwise node interactions, and thus factorize the probability of the graph adjacency matrix into the dot product of factor representations of nodes. Sequential models (You et al., 2018; Wang et al., 2018; Jin et al., 2018; Han et al., 2023) adopt a similar concept of node interactions but correlate these interactions by organizing them into a series of connection events. Additionally, some models treat the graph adjacency matrix as a parameterized random matrix and generate it by mapping a random vector through a feed-forward neural network (Simonovsky and Komodakis, 2018; De Cao and Kipf, 2018). In terms of optimization targets, many utilize log-likelihood-based objectives such as negative log-likelihood (Liu et al., 2019) or evidence lower bound objectives (Kipf and Welling, 2016), while others employ generative adversarial losses (Wang et al., 2018) or reinforcement learning losses (De Cao and Kipf, 2018). Diffusion-based graph generative models (Jo et al., 2022; Haefeli et al., 2022; Vignac et al., 2023; Jo et al., 2023; Chen et al., 2023; Cho et al., 2023; Kong et al., 2023), including this work, feature a unique data generation process compared to previous models. They map the observed graph structures and node features to a latent space through a stochastic diffusion process, whose reverse process can be learned by optimizing a variational lower bound (Vignac et al., 2023) or numerically solving a reverse stochastic differential equation (Jo et al., 2022).
Diffusion models.
The stochastic diffusion process is introduced by Sohl-Dickstein et al. (2015) for deep unsupervised learning, and its foundational connection with deep generative models is laid down by denoising diffusion probabilistic model (DDPM) (Ho et al., 2020). DDPM maps a data sample to the latent space via a Markov process that gradually applies noise to the original sample, and learns a reverse process to reproduce the sample in finite steps. The optimization and sampling processes in DDPM can be interpreted through the lens of variational inference (Sohl-Dickstein et al., 2015; Ho et al., 2020) or can be formulated as score matching with Langevin dynamics (Song et al., 2021; Song and Ermon, 2019). Both approaches are focused on diffusion processes that define the transition between normally distributed variables, which are proven effective for generating natural images. As the scope of generative tasks expands to discrete domains like text, diffusion models transitioning between discrete states emerge (Austin et al., 2021), which demonstrates that the choice of probabilistic distribution for modeling each noise state could significantly impact the learning task. This conclusion is also validated in the application of graph generation (Haefeli et al., 2022; Vignac et al., 2023). Further studies (Chen and Zhou, 2023; Zhou et al., 2023) are conducted to improve diffusion models by introducing novel diffusion processes based on probabilistic distributions that better capture the intrinsic characteristics of the generation target. Among these, the beta diffusion of Zhou et al. (2023) is chosen as the foundation of our method, due to the beta distribution’s proficiency in capturing sparsity and long-tailed characteristics in range-bounded data. These traits are commonly observed in real-world graphs (Barabási et al., 2000; Ciotti et al., 2015; Liang et al., 2023; Wang et al., 2023).
4 Experiments
We validated GBD on both molecular and non-molecular benchmarks, demonstrating its ability to generate valid graphs using various evaluation metrics.
4.1 Generic graph generation
To verify that GBD can generate graphs accurately reflecting the underlying data distribution, we first evaluate our method on a range of generic generation tasks across various datasets.
Datasets and metrics.
@@ 4.1 Generic graph generation
We consider three synthetic and real datasets of varying size and connectivity used as benchmarks in previous works: Ego-small consists of 200 small real sub-graphs from the Citeseer network dataset with $4\leq N\leq 18$, where $N$ denotes the number of nodes. Community-small consists of 100 randomly generated synthetic graphs with $12\leq N\leq 20$. Grid consists of randomly generated 100 standard 2D grid graphs with $100\leq N\leq 400$.
For a fair comparison, we follow the experimental and evaluation setting of Jo et al. (2022, 2023) with the same train/test split. We adopt maximum mean discrepancy (MMD) (Gretton et al., 2012) as our evaluation metric to compare three graph property distributions between the test graphs and the same number of generated graphs: degree (Deg.), clustering coefficient (Clus.), count of orbits with 4 nodes (Orbit), and their average score (Avg.). Further details about these metrics can be found in Appendix A.
Baselines.
We compare GBD against the following autoregressive and one-shot graph generation methods: DeepGMG (Li et al., 2018), GraphRNN (You et al., 2018), GraphAF (Shi et al., 2020), GraphDF (Luo et al., 2021), GraphVAE (Simonovsky and Komodakis, 2018), and GNF (Liu et al., 2019). We also compare GBD against several state-of-the-art diffusion-based graph generative models: EDP-GNN (Niu et al., 2020a), a score-based model for adjacency matrix, GDSS (Jo et al., 2022) and ConGress (Vignac et al., 2023), continuous diffusion models, DiGress (Vignac et al., 2023), a discrete diffusion model, and Wave-GD (Cho et al., 2023), a wavelet-based diffusion model. We describe implementation details in Appendix A.
Results.
As shown in Table 1, the proposed GBD achieves superior or comparable performance to the MMD of the sampled graphs from the training set. Compared to previous diffusion-based methods, GBD outperforms them on 7/9 MMD metrics and achieves the best average results on each dataset. In particular, GBD surpasses DiGress on all MMD metrics, demonstrating that our model is capable of modeling discrete graph data and even achieves better results. We attribute this to our model’s superior ability to capture the sparsity of the graph data distribution, which makes GBD distinctive. For the larger graph dataset Grid, GBD obtains a superior average score and other competitive MMDs compared to Wave-GD, demonstrating its potential in generating larger graphs.
Evaluation with larger sample size.

### v2

@@ C.1 Training and Sampling
C.2 Details of Concentration Modulation
Here we elaborate the concentration modulation strategies mentioned in Section 2.3.
Concentration Modulation for general graph generation
For general graph generation, we provide two strategies depending on node-level centralities, which are degree centrality and betweenness centrality, respectively. For the degree centrality of the node $u$ in an undirected graph, it can be formulated as
${C_{d}}(u)=C_{in}(u)=C_{out}(u)=\text{Deg}(u)$ (30)
where $C_{in}(u)$, $C_{out}(u)$ and $\text{Deg}(u)$ are denoted as the in-degree, out-degree, and total degree of the node $u$. For the betweenness centrality of the node $u$ in an undirected graph, it can be formulated as
${C_{b}}(u)=\sum\nolimits_{s\neq t\neq u}\frac{g_{st}(u)}{g_{st}}$ (31)
where $g_{st}$ is the total number of shortest paths from node $s$ to node $t$, $g_{st}(u)$ is the number of those paths that pass through $v$. For large graphs, exact calculation of betweenness centrality can be time-consuming, thus approximation algorithms using random sampling are often employed. In practice, we utilize the library of NetworkX Hagberg et al. (2008) to implement this.
With these two metrics to measure the centrality of the nodes, which are denoted as $C(u)$ in general, the modulated $\eta$ can be mathematically expressed as
$\eta_{u,v}=g_{A}(\max(\mathrm{C}(u),\mathrm{C}(v))),\quad\eta_{u}=g_{X}(\mathrm{C}(u)).$ (32)
where $g_{A}(\cdot)$ and $g_{X}(\cdot)$ are two assignment functions that map the node centrality to one of the predefined $\eta$ values.
Concentration Modulation for molecule graph generation
For molecule graph generation, we provide a straightforward strategy that regards the carbon atom as the most important node, as well as the bonds connected to the carbon atoms as the important edge in a molecule graph. Specifically, for various types of carbon-atom bond, we first rank the importance of bonds in the following order: carbon-carbon bonds, carbon-nitrogen bonds, carbon-oxygen bonds, carbon-sulfur bonds, and so on. Then we can assign predefined $\eta$ on different nodes and edges depending on their "importance" in a molecule graph.
C.3 Model Architechture
We leverage the graph transformer network introduced in Dwivedi & Bresson (2020); Vignac et al. (2023) cross all graph generation tasks. Each graph transformer layer consists of a graph attention module, as well as fully-connected layers and layer normalization. It employs self-attention module to update node features, then uses FiLM layers (Perez et al., 2018) to incorporate edge features and global features. Since the data we transformed falls within the range of [0, 1], we apply the sigmoid function to the output of node features and adjacency matrices to model the one-hot encoded representation of node and edge.
C.4 Schedule of Diffusion Process in GBD
Following Zhou et al. (2023), we employ a sigmoid diffusion schedule defined as $\alpha_{t}=1/(1+e^{-{c_{0}}-(c_{1}-c_{0})^{t}})$ throughout all experiments, where $c_{0}=10$ and $c_{1}=-13$
Appendix D Experimental Details
D.1 General graph generation
Datasets
@@ D.1 General graph generation
We evaluated our model using three synthetic and real datasets of varying size and connectivity, previously used as benchmarks in the literature (Cho et al., 2023; Jo et al., 2022): Ego-small (Sen et al., 2008) consists of 200 small real sub-graphs from the Citeseer network dataset with $4\leq N\leq 18$. Community-small consists of 100 randomly generated synthetic graphs with $12\leq N\leq 20$, where the graphs are constructed by two equal-sized communities, each of which is generated by the Erdös–Rényi model (Erdős et al., 1960), with $p=0.7$ and $0.05N$ inter-community edges are added with uniform probability as in previous works (Jo et al., 2022; Niu et al., 2020b). Grid consists of randomly generated 100 standard 2D grid graphs with $100\leq N\leq 400$ and the maximum number of edges per node is 4 since all nodes are arranged in a regular lattice. Planar comprises 200 synthetic planar graphs, each containing $N=64$. nodes. SBM includes 200 synthetic stochastic block model graphs, where the number of communities is randomly sampled from 2 to 5, and the number of nodes in each community is randomly sampled from 20 to 40. The probability of edges between communities is 0.3 and that of edges within communities is 0.05.
Evaluation metrics
For a fair comparison, we follow the experimental and evaluation settings of Jo et al. (2022, 2023), using the same train/test split, where 80% of the data is used as the training set and the remaining 20% as the test set. We adopt maximum mean discrepancy (MMD) as our evaluation metric to compare three graph property distributions between test graphs and the same number of generated graphs: degree (Deg.), clustering coefficient (Clus.) and count of orbits with 4 nodes (Orbit). Note that we use the Gaussian Earth Mover’s Distance (EMD) kernel to compute the MMDs following the method used in previous work (Jo et al., 2022; Cho et al., 2023). Additionally, for graphs with more complex structures like Planar and SBM, we report the eigenvalues of the graph Laplacian (Spec.) and the percentage of valid, unique, and novel graphs (V.U.N.) to measure whether the model has learned the intrinsic feature distribution and global properties of the graph. A lower value is better for all of these metrics except V.U.N. Specifically, a graph is defined as a valid planar graph if it is connected and planar. Following the statistical test introduced in  Martinkus et al. (2022), we determine that a graph is a valid SBM graph if and only if it has a community count between 2 and 5, and a node count inside each community between 20 and 40.
Implementation details
We follow the evaluation setting of  Jo et al. (2022); Cho et al. (2023) to generate graphs of the same size as the test data in each run and we report the mean and standard deviation obtained from 3 independent runs for each dataset. We report the baseline results taken from  Cho et al. (2023), except for the results of ConGress in Tables 1 and 8, which we obtained by running its corresponding open-source code. For a fair comparison, we adopt the Graph Transformer (Dwivedi & Bresson, 2020; Vignac et al., 2023) as the neural network used in GDSS+Transformer (Jo et al., 2022), DiGress (Vignac et al., 2023), and DruM (Jo et al., 2023). We set the diffusion steps to 1000 for all the diffusion models. For important hyperparameters mentioned in Sec  2.3, we usually set $S_{cale}=0.9$, $S_{hift}=0.09$. and $\eta=[10000,100,30,10]$ for the normalized degrees corresponding to the intervals falling in the interval split by $[1.0,0.8,0.4,0.1]$, respectively. In practice, we set threshold as 0.5 to quantize generated continue adjacency matrix.
D.2 2D molecule generation
Datasets
We utilize two widely-used molecular datasets as benchmarks, as described in Jo et al. (2023): QM9 (Ramakrishnan et al., 2014), consisting of 133,885 molecules with $N\leq 9$ nodes from 4 different node types and ZINC250k (Irwin et al., 2012), consisting of 249,455 molecules with $N\leq 38$ nodes from 9 node types. Molecules in both datasets have 3 edge types, namely single bond, double bond, and triple bond. Following the standard procedure in the literature (Shi et al., 2020; Luo et al., 2021; Jo et al., 2022, 2023), we kekulize the molecules using the RDKit library (Landrum et al., 2006) and remove the hydrogen atoms from the molecules in the QM9 and ZINC250k datasets.
Evaluation metrics
Following the evaluation setting of  Jo et al. (2022), we generate 10,000 molecules for each dataset and evaluate them with four metrics: the ratio of valid molecules without correction (Val.). Frechet ChemNet Distance (FCD) evaluates the chemical properties of the molecules by measuring the distance between the feature vectors of generated molecules and those in the test set using ChemNet. Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) assesses the quality of the graph structure by measuring the MMD between the generated molecular graphs and the molecular graphs from the test set. Scaffold Similarity (Scaf.) evaluates the ability to generate similar substructures by measuring the cosine similarity of the frequencies of Bemis-Murcko scaffolds (Bemis & Murcko, 1996).
Implementation details
We follow the evaluation setting of  Jo et al. (2022, 2023) to generate 10,000 molecules and evaluate graphs with test data for each dataset. We quote the baselines results from  Jo et al. (2023). For a fair comparison, we adopt the Graph Transformer (Dwivedi & Bresson, 2020; Vignac et al., 2023) as the neural network used in GDSS+Transformer (TF) (Jo et al., 2022), DiGress (Vignac et al., 2023), and DruM (Jo et al., 2023). We apply the exponential moving average (EMA) to the parameters while sampling and set the diffusion steps to 1000 for all the diffusion models. For both QM9 and ZINC250k, we encode nodes and edges to one-hot and set $S_{cale}=0.9$, $S_{hift}=0.09$. For $\eta$ modulated in molecule generation, with the help of chemical knowledge, we apply $\eta=[10000,100,100,100,30]$ to carbon-carbon bonds, carbon-nitrogen, carbon-oxygen, carbon-fluorine, and other possible bonds, respectively. For $\eta$ of nodes, we apply $\eta=[10000,100,100,30]$ on carbon atom, nitrogen atom, oxygen atom and other possible atoms, respectively. As described in Section 2.3, applying the appropriate $\eta$ for different node types and edge types can prolong the presence of related substructures during the diffusion process. In practice, we set threshold as 0.5 to quantize generated continue adjacency matrix, and the value in discrete adjacency matrix is 0 after quantizing if and only if all values in each dimension are all 0.
Complete results on 2D molecule generation
We provided additional results including Unique and Novelty on 2D molecule generation in Table 6 and Table 7.
Table 6: Additional 2D molecule generation results on QM9 dataset.
QM9 ($|N|\leq 9$)
Method Valid (%) $\uparrow$ FCD $\downarrow$ NSPDK $\downarrow$ Scaf. $\uparrow$ Uniq (%) $\uparrow$ Novelty (%) $\uparrow$
MoFlow 91.36 4.467 0.0169 0.1447 98.65 94.72
GraphAF 74.43 5.625 0.0207 0.3046 88.64 86.59
GraphDF 93.88 10.928 0.0636 0.0978 98.58 98.54
EDP-GNN 47.52 2.680 0.0046 0.3270 99.25 86.58
GDSS 95.72 2.900 0.0033 0.6983 98.46 86.27
GDSS+TF 99.68 0.737 0.0024 0.9129 - -
DiGress 98.19 0.095 0.0003 0.9353 96.67 25.58
SwinGNN 99.71 0.125 0.0003 - 96.25 17.34
GraphARM 90.25 1.22 0.0020 - 95.62 70.39
EDGE 99.10 0.458 - 0.763 100.0 -
GruM 99.69 0.108 0.0002 0.9449 96.90 24.15
GBD 99.88 0.093 0.0002 0.9510 97.12 26.32
Table 7: Additional 2D molecule generation results on ZINC250k dataset.
ZINC250k ($|N|\leq 38$)
Method Valid (%) $\uparrow$ FCD $\downarrow$ NSPDK $\downarrow$ Scaf. $\uparrow$ Uniq (%) $\uparrow$ Novelty (%) $\uparrow$
MoFlow 63.11 20.931 0.0455 0.0133 99.99 100.0
GraphAF 68.47 16.023 0.0442 0.0672 98.64 99.99
GraphDF 90.61 33.546 0.1770 0.0000 99.63 100.0
EDP-GNN 82.97 16.737 0.0485 0.0000 99.79 100.0
GDSS 97.01 14.656 0.0195 0.0467 99.64 100.0
GDSS+TF 96.04 5.556 0.0326 0.3205 - -
DiGress 94.99 3.482 0.0021 0.4163 99.97 99.99
SwinGNN 81.72 5.920 0.006 - 99.98 99.91
GraphARM 88.23 16.26 0.055 - 99.46 100.0
GruM 98.65 2.257 0.0015 0.5299 99.97 99.98
GBD 97.87 2.248 0.0018 0.5042 99.97 99.99
D.3 Computing resources
For all experiments, we utilized the PyTorch (Paszke et al., 2019) framework to implement GBD and trained the model with NVIDIA GeForce RTX 4090 and RTX A5000 GPUs.
Appendix E Additional Experimental Results
E.1 Evaluation with larger sample size